OpenStudy (anonymous):

Suppose there are three white marbles and three black marbles in a bag and you want to remove two marbles. What is the probability that you will select a white marble and then a black marble? Express your answer as a percent.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yeah 50% because there is an equal amount of marbles

OpenStudy (anonymous):

ok, just wanted to be sure i was right

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

:)

OpenStudy (phi):

The problem says you want to remove two marbles. So you do not replace the marble back in the bag. This changes things. probability of getting a white marble on the first pick is # of white / total number = 3/6 =1/2 now there are 5 marbles left in the bag: 2 white and 3 black. Probability of drawing a black marble is 3/5 The probability of getting a white followed by a black is (1/2)*(3/5) = 3/10 or 30%